OpenStudy (anonymous):

1 of 24 (3 pt) Which expression represents an unknown number added to 13? A. 13 + t B. 13t C. 13 – t D. 13 ÷ t

OpenStudy (anonymous):

2. (3 pt) Which expression represents 76 decreased by an unknown number? A.76/g B.76-g C.76+g D.76g

OpenStudy (anonymous):

3. (3 pt) Which equation best represents this situation? The number 98 increased by an unknown number is equal to 120. A. 98 – x = 120 B. 98 + x = 120 C. 98 * x = 120 D. 98 ÷ x = 120

OpenStudy (igreen):

#1 "added" Which one shows addition?

OpenStudy (igreen):

By the way, only one question per post is allowed.

OpenStudy (anonymous):

4. (3 pt) Which equation best represents this situation? The number 80 minus an unknown number is equal to 46. A. 80 * h = 46 B. 80 + h = 46 C. 80 + h = 46 D. 80 - h = 46

OpenStudy (anonymous):

@iGreen it is a

OpenStudy (igreen):

Yes, A is correct.